Presure at the bottom of a tank is 3P, where P is the atmospheric pressure. If the water is drawn out till the level of water is lowered to one fifth, then what is the pressure at the bottom of the tank?

Grade:

2 Answers

Sanjeev Malik IIT BHU
12 years ago

given   dgh=2p   (d=density of water)

as h reduced to h/5, so we need dgh/5=  2p/5

so pressure at bottom= p + 2p/5

= 7p/5

 

 

click "yes" below to approve the answer....

Shambhavi Singh
8 years ago
P + dgh= 3P (d is density here)
dgh=2P

It is given that the height is lowered ‘by’ 1/5 of the original height
Therefore, 
4dgh/5=2P*4/5 (multiplying both sides by h – h/5=4/5h)
P+8P/5=13P/5